Two years ago, Lolita was thrice as old as Miraflor. Four years hence Lolita will be twice as old as Miraflor. Find their ages now.
Answer by stanbon(75887)

Two years ago, Lolita was thrice as old as Miraflor. Four years hence Lolita will be twice as old as Miraflor. Find their ages now.
----------
(L-2) = 3(M-2)
(L+4) = 2(M+4)
------
Rearrange::
L - 3M = -4
L - 2M = 4
------------
Subtract and solve for "M":
M = 8 yrs (Miraflor's age now)
---
Solve for "L":
L = 3M - 4
L = 20 yrs
========
Cheers,
Stan H.
----------